OpenStudy (anonymous):

What is the equation, in point-slope form, of the line that has a slope of 6 and passes through the point (–1, –8)?

OpenStudy (nurali):

y-y1=m(x-x1) (x1,y1)=(-1,-8) m=6 y+8=6(x+1) y+8=6x+6 y=6x-2

OpenStudy (nurali):

Welcome to OpenStudy :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct ). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)